Given that C = $2,000 + 0.60YD, if the level of disposable income is $1,000, the level of saving is $400. $600. -$600. Saving will be negative when consumption exceeds disposable income. Use the given formula where C=consumption and Y=the level of disposable income: C = $1,000 + .6($1,000) = $1,600. So when consumption of $1,600 is more than disposable income of $1,000, the difference is a decrease in saving of -$600. -$1600.
Answer #1

-$1600
(C = 2,000 + 0.60YD = 2,000 + 0.60(1000) = 2000 + 600 = 2600
S = Y - C = 1000 - 2600 = -1600

So, savings decrease by 1600.)